### v1

# Unveiling the Role of ChatGPT in Software Development: Insights from Developer-ChatGPT Interactions on GitHub

Abstract.
The advent of Large Language Models (LLMs) has introduced a new paradigm in software engineering, with generative AI tools like ChatGPT gaining widespread adoption among developers. While ChatGPT’s potential has been extensively discussed, there is limited empirical evidence exploring its real-world usage by developers. This study bridges this gap by conducting a large-scale empirical analysis of ChatGPT-assisted development activities, leveraging a curated dataset, DevChat, comprising 2,547 unique shared ChatGPT links collected from GitHub between May 2023 and June 2024. Our study examines the characteristics of ChatGPT’s usage on GitHub (including the tendency, prompt turns distribution, and link descriptions) and identifies five categories of developers’ purposes for sharing developer-ChatGPT conversations during software development. Additionally, we analyzed the development-related activities where developers shared ChatGPT links to facilitate their workflows. We then established a mapping framework among data sources, activities, and SE tasks associated with these shared ChatGPT links. Our study offers a comprehensive view of ChatGPT’s application in real-world software development scenarios and provides a foundation for its future integration into software development workflows.
1. Introduction
The emergence of Large Language Models (LLMs) has fundamentally reshaped the landscape of Software Engineering (SE) practices (Hou et al., 2023). They introduce a novel paradigm for software development, offering capabilities that can potentially revolutionize the field. As one of the most typical and representative LLM-based tools, ChatGPT (OpenAI, ChatGPT, 2025), launched by OpenAI in November 2022, has sparked widespread interest within both academia and industry, leading to numerous studies exploring its applicability in software development, such as code generation and refinement (Sobania et al., 2023; Guo et al., 2024; Li et al., 2023), code summarization and comment generation (Sun et al., 2025; Su and McMillan, 2024), and program repair (Zhang et al., 2023).
Notably, in May 2023, OpenAI released a new feature allowing users to share links within their ChatGPT interactions (ChatGPT Shared Links FAQ, 2025). This feature promotes the sharing of knowledge generated by ChatGPT and enhances team collaboration through shared developer-ChatGPT interactions. As a result, this feature has further boosted ChatGPT’s popularity among developers, surpassing other LLMs in developers’ communities. As evidenced by our trial search, before selecting ChatGPT as our target LLM, we conducted a preliminary search on GitHub for other LLMs that support shared conversation links, such as Gemini (Google, Gemini, 2025) , and found significantly fewer shared links. Consequently, we decided to use ChatGPT as the focus of this study, which is the most popular and widely used LLM on GitHub.
Recently, Xiao et al. (Xiao et al., 2024) shared a dataset of developer-ChatGPT conversations used on GitHub, collected from July to October 2023, which ignited interest in studying ChatGPT’s practical use. However, due to its limited size and temporal coverage, the dataset is not sufficient to support our comprehensive understanding of the real-world prevalence and usability of ChatGPT in software development. More specifically, we were unable to obtain some important observations and insights, such as the characteristics of ChatGPT’s usage by developers, their purposes for sharing ChatGPT links, the integration of ChatGPT into developer workflows, and the task-specific scenarios of its engagement. The integration of ChatGPT into software development workflows is still in its early stages, but it has already shown great promise in improving developer productivity and code quality (Vaillant et al., 2024). Therefore, our goal is to investigate how developers utilize ChatGPT in their daily routines to better anticipate its broader impact on the SE field.
To address these gaps and achieve our goal, we conducted a large-scale empirical investigation into shared ChatGPT links, offering insights into LLM-assisted software development. Specifically, we first curated DevChat, a large-scale dataset that encompasses ChatGPT usage from May 2023 to June 17, 2024. DevChat contains 2,547 unique shared ChatGPT links related to a variety of development activities across five main development-related data sources on GitHub, i.e., Code, Issues, Commits, Pull Requests, and Discussions. Leveraging DevChat, we conducted an in-depth analysis of how developers utilize ChatGPT during software development.
Our analysis reveals several key findings: Developers predominantly share ChatGPT links in Code (43.4%) and Commits (32.3%), and their enthusiasm for sharing ChatGPT links reached a peak in August 2023. Developers primarily employ ChatGPT for short, task-specific prompts, with two- or three-turn prompts being the most frequent. Contextual descriptions are prevalent in most shared ChatGPT links, particularly in Commits (99%), but are often absent in Code. In contrast, in Issues and Pull Requests, the absence of descriptions suggests an assumption of inherent content clarity. The developers’ purposes for sharing ChatGPT links on GitHub were identified and classified into five categories, with Task Delegation being the dominant purpose, primarily to automate repetitive work. Moreover, we categorized the development-related activities within the shared ChatGPT conversations into two primary groups: Development Activities and Supporting Activities, with Software Development, Software Maintenance and Evolution emerged as the dominant activities. Furthermore, we established the mappings among the data sources, activities, and tasks of shared developer-ChatGPT conversations. These mappings provide a comprehensive overview of 39 specific SE tasks, with Code Generation & Completion and Code Modification & Optimization emerging as the most prevalent. These findings provide crucial empirical evidence for guiding future research and the development of AI-assisted tools, contributing to a deeper understanding of the evolving AI-empowered development paradigm.
The main contributions of our study encompass:
We curated DevChat, a large-scale dataset of developer-ChatGPT interactions relevant to software development on GitHub. Following deduplication and manual review, the cleaned dataset is available in our dataset (Li et al., 2025). This dataset enables future research into the practical applications of ChatGPT in software development.
We analyzed the characteristics of ChatGPT’s usage on GitHub (including the tendency, prompt turns distribution, and link descriptions). Additionally, we identified and categorized developers’ purposes for sharing developer-ChatGPT conversations during development in practice.
We examined the development-related activities for which developers shared ChatGPT links to support their workflows. Furthermore, we established a mapping relationship among data sources, activities, and SE tasks associated with these shared links, providing a comprehensive view of ChatGPT’s role in practical development scenarios.
The remainder of this paper is organized as follows: Section 2 introduces related studies of this work. Section 3 elaborates on the process of our study design and the research questions. Section 4 describes the study results and our findings, and the discussions and implications of this study are presented in Section 5. Section 6 examines the threats to validity. Section 7 summarizes this study and outlines the future work.
2. Related Work
2.1. Large Language Models for Software Engineering
In recent years, groundbreaking advancements in Natural Language Processing (NLP) have fueled a meteoric rise in the performance and adoption of Large Language Models (LLMs). The breakthroughs allow LLMs to scale up the model sizes and significantly enhance their capacity when the parameter scale surpasses a certain threshold (Hou et al., 2023). As a result, LLMs have become increasingly powerful and versatile, driving innovation and efficiency across various fields and applications. These LLMs, such as OpenAI’s GPT series models (e.g., GPT-1 GPT-4o), Google’s Gemini (Google, Gemini, 2025), Meta’s Llama (Meta, Llama, 2025), and Anthropic’s Claude (Anthropic, Claude, 2025), are built on the Transformer architecture (Vaswani et al., 2017) to understand and generate text through the training on extensive datasets comprising diverse linguistic patterns and contexts.
LLMs are extensively utilized throughout various phases of the Software Development Life Cycle (SDLC), including requirements engineering, software design, software development, quality assurance, maintenance, and management (Hou et al., 2023). The most predominant phase is software development, and code generation and program repair being the most prevalent tasks for LLMs (Hou et al., 2023). For example, as a code generation tool empowered by LLM, GitHub Copilot (GitHub, Copilot, 2025) received much attention from software developers after its release. Vaithilingam et al. (Vaithilingam et al., 2022) conducted a user study with 24 participants on the usability of GitHub Copilot. They found that most participants preferred using Copilot, as Copilot provided a useful starting point for programming tasks and saved them the effort of searching code snippets online. Liang et al. (Liang et al., 2024) performed an exploratory study to investigate the usability of LLM-assisted programming tools, such as GitHub Copilot, with 410 programmers. Their results indicate that developers are motivated to use AI programming assistants (e.g., GitHub Copilot) because these tools help reduce keystrokes, speed up tasks, and recall syntax, but there is still a gap between developers’ needs and the tools’ output. Jin et al. (Jin et al., 2023) introduced an end-to-end program repair framework, InferFix, to fix critical security and performance bugs in Java and C#. By combining a static analyzer with a fine-tuned LLM, InferFix shows better performance than LLM baselines on patch generation in Java and C#. Moreover, a recent study of Poldrack et al. (Poldrack et al., 2023) reported the experimental results of AI-assisted coding with GPT4, including performing required coding tasks, refactoring, and improving the quality of existing code. Their results show that although AI coding tools are powerful, human involvement is still necessary to ensure the validity and accuracy of the results.
While many previous user studies have explored the potential and characteristics of LLM-assisted programming, it remains unclear which development-related activities pertain to leveraging LLMs and the characteristics of their usages. Therefore, our study further extends and validates LLM’s (especially ChatGPT) usage during software development.
2.2. ChatGPT in Software Development
Researchers prefer ChatGPT over other LLMs and LLM-based applications due to its computational efficiency, adaptability to various tasks, and potential cost-effectiveness (Hou et al., 2023). As a result, the use of ChatGPT in software development encompasses a wide spectrum of applications. It ranges from generating code snippets and debugging to providing documentation, and ChatGPT proves to be a versatile tool for various stages of the development process. Hao et al. (Hao et al., 2024) conducted an empirical study to investigate the role of ChatGPT in collaborative coding. They analyzed developers’ shared conversations with ChatGPT on GitHub pull requests and issues. Their results show that developers seek ChatGPT’s assistance across 16 types of SE tasks and frequently engage with ChatGPT via multi-turn conversations; developers leverage shared ChatGPT conversations to facilitate various contributions (e.g., as code reviewers). Sobania et al. (Sobania et al., 2023) evaluated and analyzed the automatic bug-fixing performance of ChatGPT. Their findings show that ChatGPT’s program repair performance is competitive with the results achieved with common deep learning approaches like CoCoNut (Lutellier et al., 2020) and Codex (Prenner et al., 2022), and notably better than the results reported for the standard program repair approaches.
Besides, certain concerns about the correctness of ChatGPT-generated information have been raised. Kabir et al. (Kabir et al., 2024) empirically investigated the characteristics of ChatGPT answers to 517 programming questions in Stack Overflow alongside manual analysis, linguistic analysis, and user study. Their findings indicate that ChatGPT produces incorrect answers in more than 50% of the cases, and manual validation identified a large number of conceptual and logical errors in the ChatGPT answers. Li et al. (Li et al., 2023) proposed a new paradigm for finding failure-inducing test cases through leveraging ChatGPT. They found that ChatGPT initially struggled with pinpointing the buggy code, especially when program versions had subtle differences. To address this weakness of ChatGPT, the authors designed a novel approach by leveraging ChatGPT’s ability to infer expected behavior from erroneous programs and amplify the subtle code differences. Their experimental results show that the approach can significantly improve the accuracy of identifying fault-inducing test cases. Vaillant et al. (Vaillant et al., 2024) surveyed 207 developers to investigate the impact of ChatGPT on software quality, productivity, and job satisfaction. Their results show that around 70% of developers perceive ChatGPT as improving productivity and job satisfaction. Aguiar et al. (Aguiar et al., 2024) conducted an empirical study to analyze developers’ use of ChatGPT in multi-language software development. They found that 70% of the developer-ChatGPT conversations are seeking coding support regarding multiple programming languages, while 57% of developers used ChatGPT as a tool to generate code in multiple languages.
Despite ChatGPT’s popularity with developers, a comprehensive analysis of its responses within software development contexts remains absent. This study aims to bridge this gap by empirically investigating ChatGPT’s usage on GitHub. Compared to the above studies, our work conducted an in-depth investigation of practical ChatGPT usage in real-world scenarios based on our established large and clean dataset (i.e., DevChat (Li et al., 2025)).
3. Study Design
In this section, we present the Research Questions (RQs) (Section 3.1) used to achieve the goal stated in the Introduction section (Section 1) and provide an overview of the DevChat dataset curation process containing data collection (Section 3.2), data cleaning (Section 3.3), data labeling (Section 3.4), and data analysis (Section 3.5). Figure 1 shows an overview of the DevChat dataset curation process.
3.1. Research Questions
Rationale: RQ1 aims to understand how developers interact with ChatGPT, such as the frequency and depth of their prompts, as well as how and where they share developer-ChatGPT conversations. By examining trends in ChatGPT usage, the associated prompt turns distribution, and the descriptions accompanying shared ChatGPT links, this RQ can help to disclose the characteristics of ChatGPT’s usage throughout the development process.
Rationale: By exploring the diverse purposes for which developers utilize ChatGPT, RQ2 seeks to provide nuanced insights into the practical utility, potential benefits, and limitations of ChatGPT in development, thereby laying the groundwork for more effective integration of Generative AI (GenAI) tools in software development workflows.
Rationale: RQ3 is designed to map out a comprehensive landscape of ChatGPT used during software development. By identifying and categorizing various development-related activities where developers share developer-ChatGPT conversation links, this RQ quantitatively examines the distribution of ChatGPT links, offering insights into ChatGPT’s practical applications and impact on the development process.
Rationale: RQ4 delves into analyzing the specific tasks of the shared ChatGPT conversations engaged in SE activities, for example, by analyzing task-specific interactions (e.g., code-level and design-level tasks). This RQ seeks to systematically categorize the types of ChatGPT-assisted tasks in distinct SE activities, and derive actionable insights for optimizing ChatGPT’s utility and efficiency in SE practices.
3.2. Data Collection
We curated our DevChat dataset (Li et al., 2025) by collecting conversation records of developers interacting with ChatGPT across five primary data sources on GitHub, specifically targeting Code, Commits, Pull Requests, Issues, and Discussions, as these five sources encompass rich data and effectively capture developer interactions. Our data collection employed a hybrid approach, leveraging both GitHub’s REST API (Github REST API documentation, 2022) and Web crawling techniques. For data accessible via direct queries, we constructed RESTful query strings tailored to the API’s specifications, enabling us to search for shared ChatGPT conversation records and retrieve the relevant information. For data where direct access was not feasible, we built a Web crawler to capture the required information.
Since OpenAI introduced the ChatGPT conversation-sharing feature in May 2023 using a standardized URL format of “chat.openai.com/share/{conversation_id}”, our data collection on GitHub was confined to the period from May 2023 to the collection cutoff date (June 17, 2024) when we started this work. Hence, the targeted search string is settled as “chat.openai.com/share” to retrieve the shared ChatGPT conversation records.
Due to GitHub’s REST API limit of a maximum of 1,000 search results per query (REST API endpoints for search, 2025), it was not feasible to retrieve all searched entries directly from each data source. To create the most comprehensive dataset, we employed several strategies to maximize the volume of search results, adapting our approach to each of the five GitHub data sources given variations in search results (e.g., for slicing a certain time interval and programming languages). We detail our approach for collecting data from each source below:
Code: The majority of relevant results were from the Code repository and most of them were invalid (e.g., incomplete URLs). Therefore, we conducted language-based searches, constructing query strings for the top 50 programming languages on GitHub (Github Language Stats, 2025) (e.g., Python, Java, and Go), along with Markdown and HTML, which are commonly used by developers. The results from each language-specific search were then consolidated into a unified dataset.
Commits: Regarding the ChatGPT conversation records from the Commits repository, no additional filtering was necessary, as the total number of search results was below 1,000 at the time of data collection.
Issues and Pull Requests: As GitHub’s REST API treats both Issues and Pull Requests under a single “issue” endpoint, and the total number of search results exceeded 1,000, we employed a two-step filtering approach. First, we separated issues and pull requests using the “is:pr” and “is:issue” qualifiers, respectively. Second, we implemented a time-slicing strategy, dividing the search period from May 2023 (the release date of ChatGPT’s sharing feature) to June 17, 2024 into monthly intervals. For each monthly slice, we used the “created” qualifier to retrieve issues and pull requests containing the target URL substrings. The results of all monthly slices were then combined to form a complete dataset.
Discussions: Due to the absence of GitHub’s REST API for the Discussions repository, we developed a Web crawler to extract ChatGPT conversation records from the search results on GitHub Discussions.
We accessed each conversation URL to collect developer-ChatGPT interactions in a prompt-response format, storing the collected data in JSON files. As shown in Table 1, the initial data collection results are presented in the Initial Collected Data column.
Source | Initial Collected Data | Invalid URLs | Non-English | Data Cleaning | Data Labeling |
Code | 3,434 | 77 | 1,518 | 1,839 | 1,105 |
Commits | 977 | 12 | 29 | 936 | 823 |
Issues | 2,352 | 1,321 | 616 | 415 | 295 |
Pull Requests | 899 | 382 | 216 | 301 | 266 |
Discussions | 176 | 77 | 0 | 99 | 58 |
Total | 7,838 | 1,869 | 2,379 | 3,590 | 2,547 |
3.3. Data Cleaning
After collecting accessible data from GitHub, we performed preliminary data cleaning to eliminate invalid search results based on the following Exclusion (E) criteria. Statistical results for the exclusions are presented in the Invalid URLs and Non-English columns of Table 1.
E1. Invalid URLs: Those invalid ChatGPT conversation URLs could be retrieved in the initial collected data for various reasons: 1) Canceled Sharing: the users who shared the conversation records may have subsequently canceled the sharing, rendering the URLs inaccessible; 2) Processed URLs: the retrieval results may include processed ChatGPT conversation URLs, resulting in entries like “chat.openai.com/share/xxx [invalid URL hashcode]”, which do not correspond to valid conversation links; 3) Invalid URL formant: some users shared ChatGPT conversation links with incorrect formats, such as “chat.openai.com/share/e/5ec5788d-8151-4c7e-ab1e-68d84a80b79f”, in which the inclusion of “e/” deviates from the standard format. Even after removing “e/”, the links remained inaccessible. Due to link disruption, we treated those instances as invalid links, as their original content could not be retrieved.
E2. Non-English text: The search results contain non-English text (except emoji), shown in the Non-English column of Table 1. Developers might use ChatGPT for tasks such as document internationalization; therefore, we excluded these non-English search results.
3.4. Data Labeling
To streamline subsequent data labeling, we extracted relevant fields (see Section 3.5) from the JSON files obtained in the previous step. These fields were primarily selected to facilitate the data labeling process and were stored in MS Excel sheets for annotation.
After the initial data cleaning, we further manually checked and labeled the initial cleaned data during the data labeling process. Meanwhile, we also needed to remove duplicate data. For example, in many cases, developers tend to create or modify multiple files within a single commit. However, the corresponding commit messages often only provide a single shared ChatGPT link, leading to multiple GitHub links for different files being associated with the same shared developer-ChatGPT conversation. To address this, it is necessary to retain only one data entity related to the shared ChatGPT links and remove redundant data entities. Our approach to data preprocessing, specifically the removal of duplicate data entities, slightly differs from that of other studies (e.g., studies based on the DevGPT dataset (Xiao et al., 2024)). To achieve our goal of understanding how developers utilize ChatGPT during the development process, we eliminated duplicate ChatGPT links (data entities sharing a single ChatGPT URL across different GitHub links). This ensures the robustness of our statistical results and data analysis.
In terms of data labeling, all the authors discussed together to specify the data labeling task, and then the first, third, and fourth authors conducted a pilot data labeling before the formal data labeling. We used Cohen’s Kappa coefficient (Cohen, 1960) to measure the reliability and consistency of the labeled data. For the pilot data labeling, the first author reached an agreement rate of 0.79 and 0.65 with the third and fourth authors, respectively, which represents that a substantial agreement was achieved between the authors (Landis and Koch, 1977). All authors reached a consensus through discussions when encountering disagreements on the labels. The formal data labeling adhered to the methodology established during the pilot phase. Any discrepancies were resolved through collaborative discussions among all the authors.
3.5. Data Analysis
Our DevChat dataset was curated from five sources on GitHub, as detailed in Table 1. For each data source, we provided distinct metadata in JSON files to support the data analysis. Each JSON file comprises data entities with the common data items, including basic GitHub repository information (e.g., Type, URL, Author, Title, RepoName), timestamps (e.g., AuthorAt, CreatedAt), and the extracted ChatGPT data item (i.e., ChatgptSharing). Note that, ChatgptSharing consists of detailed information related to shared ChatGPT links, such as URL of shared ChatGPT conversations, timestamps of the URLs (i.e., DateOfConversation, DateOfAccess), titles generated by ChatGPT (i.e., Title), detailed statistics of the prompts and answers (i.e., NumberOfPrompts, TokenOfPrompts, TokenOfAnswers), and the model version of ChatGPT used (i.e., Model).
In this work, we examine how developers leverage ChatGPT throughout the software development process. For the data analysis, we utilized statistical methods to address the characteristics (RQ1) associated with shared ChatGPT links. To examine and classify the purposes (RQ2), activities (RQ3), and SE tasks (RQ4) of employing ChatGPT on software development, we adopted the Constant Comparison method (Charmaz, 2014; Stol et al., 2016). According to Charmaz et al. (Charmaz, 2014), the Constant Comparison process involves three key steps. In the first step, initial coding, three authors (the first, third, and fourth authors) independently reviewed the content of shared ChatGPT conversations. This was followed by focused coding, where the three authors selected the most frequent codes and categorized the data accordingly, with the results subsequently reviewed by the first author. For instance, we consolidated codes such as “Dataset Split”, “Data Merge”, and “Data Cleaning” into “Data Handling”, and merged “Program Comprehension” into “Code Explanation and Understanding”. Third, we conducted theoretical coding to establish relationships between the identified codes. To minimize personal bias, disagreements in coding results were resolved through collaborative, multi-turn discussions among all the authors, ensuring an objective interpretation of the data. Note that, we extracted and stored relevant data entities from shared ChatGPT links in MS Excel files (see the dataset at (Li et al., 2025)). Our dataset also lends itself to further analysis of other topics, such as prompts, responses, and relevant interactions.
4. Results
In this section, we present the results of our data analysis and address the RQs outlined in Section 3.1.
4.1. RQ1: Characteristics of ChatGPT Usages
To answer RQ1, we analyzed the characteristics of ChatGPT usage based on our curated dataset, including the distribution of shared ChatGPT links on GitHub, the associated prompts, and link descriptions.
(1) Data distribution of shared ChatGPT links on GitHub
Figure 2 illustrates the distribution of 2,547 shared ChatGPT links across five data sources on GitHub (see the Donut diagram) and the monthly usage of ChatGPT by developers during software development on GitHub (see the Line diagram). This figure provides insights into the contexts where developers most commonly interact with ChatGPT within their workflow.
Notably, Code emerges as the dominant share of interactions, comprising 43.4% (1,105) of all shared URLs. This significant share underscores ChatGPT’s primary role in supporting code-related tasks, such as generating code snippets or troubleshooting issues within the codebase. Commits follow as the second largest share at 32.3% (823), indicating that developers also turn to ChatGPT for support in committing changes, potentially refining, validating, or documenting code changes before integration. The remaining sources show comparatively lower levels of ChatGPT engagement. Issues account for 11.6% (295), implying moderate use of ChatGPT for tracking and resolving problems. Pull Requests contribute 10.4% (266), suggesting that developers sometimes employ ChatGPT during code review and merging processes. Lastly, Discussions make up a mere 2.3% (58), reflecting limited reliance on ChatGPT for open-ended or collaborative discussions. The distribution of shared ChatGPT links reveals a strong preference for ChatGPT in focused and procedural work like coding and committing. In contrast, its use in more interactive and exploratory exchanges is comparatively lower, such as discussions.
The Line diagram in Figure 2 illustrates the monthly trend of shared ChatGPT links across the five GitHub data sources from January 2023 to May 2024 (note that a few collected ChatGPT URLs were created before the release of ChatGPT’s sharing feature, i.e., May 2023). Following OpenAI’s release of the shared feature, a notable surge in usage occurred, peaking around August 2023, demonstrating significant initial adoption. While usage fluctuates across sources, Code and Commits consistently exhibit high levels of engagement, reinforcing their importance in developers’ daily workflows. In contrast, Discussions consistently exhibits lower usage, likely due to the less popularity of open-ended conversations on GitHub, and developers tend to prioritize concise, task-oriented interactions over more extensive, exploratory conversations.
From another perspective, Figure 3 illustrates the cumulative adoption of ChatGPT by developers. The steady growth of ChatGPT’s usage across the five sources highlights the platform’s popularity and importance in the software development community. The lines for Code and Commits consistently remain the most prominent, which aligns with GitHub’s primary role as a code repository and a platform for version control. The lines for Issues and Pull Requests show a steady increase, reflecting that collaborative development and knowledge sharing are prevalent on GitHub. The Discussions line remains relatively low compared to the others, suggesting that discussions might not be as heavily utilized on GitHub as other activities.
Overall, this trend reflects ChatGPT’s emerging role in software development, especially in collaborative and code-centric work on GitHub.
(2) Prompt turns distribution of shared ChatGPT links on GitHub
Our DevChat dataset also contains the prompt data of developer-ChatGPT interactions. Figure 4 provides a box plot visualization of the prompt turns distribution across five GitHub sources, after filtering outliers. The associated statistics are summarized in Table 2, which includes the first (Q1), second (Q2, a.k.a Median), and third (Q3) quartiles of the distributions.
The high average number of prompt turns in shared ChatGPT conversations observed in Code indicates the inherently iterative and complex nature of missions like code generation, debugging, and explanation. Moreover, the prompt turns of developer-ChatGPT interactions in Code are higher than those found in other data sources. The distribution of prompt turns in Commits, Issues, Pull Requests, and Discussions indicates developers’ usage of ChatGPT for short and task-focused interactions (e.g., concept interpretation, code modification).
Overall, multi-turn prompts within these interactions highlight ChatGPT’s versatility, supporting not only concise tasks but also deeper discussions and collaborative problem-solving.
Source | Count | Q1 | Q2 | Q3 | Min/Max |
Code | 1,105 | 1 | 3 | 7 | 1 / 302 |
Commits | 823 | 1 | 1 | 2 | 1 / 89 |
Issues | 295 | 1 | 2 | 4 | 1 / 109 |
Pull Requests | 266 | 1 | 2 | 4 | 1 / 167 |
Discussions | 58 | 1 | 2 | 4 | 1 / 82 |
(3) Description of shared ChatGPT links on GitHub
To investigate the descriptions of shared ChatGPT links, we examined the description distribution of the shared ChatGPT links across the five GitHub sources. We identified two primary categories when developers share ChatGPT links: sharing ChatGPT links with or without a description. Figure 5 presents an overview of the two categories regarding developer-ChatGPT interactions.
Figure 5 reveals a strong tendency to include contextual description when sharing ChatGPT links, especially in collaborations like Commits (99.39%), Discussions (91.38%), and Pull Requests (89.72%). In contrast, the practices of attaching descriptions to shared ChatGPT links in Code vary, 70.95% and 29.05% of the shared ChatGPT links with and without a description, respectively. More specifically, nearly half (47.33%) of the shared ChatGPT links in code comments provide a contextual description, while 23.17% lack contextual information. Shared links lacking descriptions are minimal in README files (2.62%) and code snippets (0.81%). This indicates that developers may rely on the link content to interpret or assume that the recipient has the necessary background to understand the link content.
Overall, 82.80% of shared ChatGPT links contain descriptions, whereas 17.20% of them do not offer any contextual information. The findings reflect that developers generally prefer to provide descriptive information when sharing ChatGPT links, possibly to ensure clarity and facilitate collaborations in development with the shared developer-ChatGPT interactions.
4.2. RQ2: Developers’ Purposes for Sharing ChatGPT Conversations
To answer RQ2, we employed the Constant Comparison method (Charmaz, 2014; Stol et al., 2016) to analyze and categorize the developers’ purposes for sharing ChatGPT links during software development. We finally got five primary categories of developers’ purposes (i.e., Task Delegation, Problem Resolution, Knowledge Acquisition, Solution Recommendation, and Concept Interpretation), and we analyzed their prevalence across the five sources. Figure 6 presents the distribution of the five purposes across five sources. Each circle is scaled to the total number of shared ChatGPT links for different sources, and the segments within each circle show the relative contribution of each purpose, making cross-category comparisons intuitive. Detailed results are presented below.
(1) Task Delegation encompasses the use of ChatGPT to automate or execute specific development tasks. Task delegation emerges as the predominant purpose, particularly in Commits (85.66%) and Pull Requests (88.35%), reflecting ChatGPT’s role in automating repetitive tasks. For instance, developers frequently employ ChatGPT to generate boilerplate code and README files. The high prevalence of task delegation in Code (75.02%) further underscores the utility of task delegation in accelerating code creation. However, the lower adoption of task delegation in Issues (61.02%) and Discussions (50.00%) suggests that collaborative or context-rich tasks may demand human inspection, limiting the automation potential in these sources.
(2) Problem Resolution involves debugging code, fixing errors, or explaining error messages. Problem resolution peaks in Commits (11.42%) and Issues (8.47%), aligning with scenarios where developers address problems when merging code or triaging issues. The minimal presence of problem resolution in Pull Requests (1.13%) implies that ChatGPT may not be as commonly relied upon for resolving complex or collaborative challenges. However, ChatGPT’s utility in problem resolution underscores its potential as an efficient tool to identify and fix errors during development.
(3) Knowledge Acquisition refers to the proactive process that developers seek for clarifications, differences, and domain-specific knowledge during development to contribute effectively to their projects. Knowledge acquisition is relatively prominent in Discussions (20.69%) and Issues (12.88%), where developers seek clarifications or compare technical approaches relevant to development. While ChatGPT’s role in knowledge acquisition is not as dominant in Commits (0.61%) and Pull Requests (4.51%), ChatGPT’s contribution to the educational aspects of development is noteworthy. By serving as a knowledge repository and a search engine, ChatGPT helps developers quickly access relevant details and broaden their understanding of various knowledge points.
(4) Solution Recommendation covers requests for solutions, recommendations, suggestions, etc, to guide implementation or resolve issues efficiently. Solution Recommendation is comparatively more frequent in Discussions (17.24%) and Issues (12.54%), indicating ChatGPT’s value in collaborative problem-solving. Developers appear to use ChatGPT to generate actionable suggestions during ideation, such as recommending design patterns, best practices, and decision-making for making sound technical decisions. The lower prevalence of this task in Code (2.26%) and Pull Requests (2.63%) implies that GitHub’s usage conventions that developers focus on execution (e.g., implement predefined and specific tasks like bug fixing) rather than exploration (e.g., recommend and try new possibilities like designing algorithms) in these two sources on GitHub (Wessel et al., 2023).
(5) Concept Interpretation: Developers usually seek ChatGPT’s help to understand or interpret specific concepts, algorithms, design principles, and other software development topics. Concept interpretation is the least adopted category overall, with modest usage in Issues (5.08%) and Discussions (5.17%). These requests typically highlight ChatGPT’s role as a development assistant. It reflects ChatGPT’s strength in facilitating concept comprehension, making ChatGPT an invaluable tool for developers seeking quick explanations or insights on unfamiliar concepts.
4.3. RQ3: Development-Related Activities Involved in Shared ChatGPT Conversations
To address RQ3, which focuses on the development-related activities where developers share ChatGPT conversations, we analyzed the distribution of shared ChatGPT links across the five data sources. This result reveals the breadth of developer interactions potentially facilitated by ChatGPT, as illustrated in Figures 7, 8, and 9.
Although there are various ways to classify software development activities, we considered and adapted the classification of software development activities proposed in SWEBOK (Washizaki et al., 2024) and previous studies (Hou et al., 2023; Fan et al., 2023), which are widely recognized in the field of software engineering. Following this, we categorized the relevant activities involving shared developer-ChatGPT conversations, as shown in Figure 7. The development-related activities were then grouped into two primary categories: Development Activities and Supporting Activities.
Development Activities encompass seven conventional phases of the software development life cycle (SDLC), that is, Requirements Analysis, Software Design, Software Development, Software Testing, Software Deployment, Software Maintenance and Evolution, and Software Management.
Supporting Activities comprise two novel AI-assisted activities that facilitate and support the software development process, i.e., Data Handling and Information Seeking.
To facilitate observation and comparison, we use two scales (0~1200 and 0~180) in Figure 8, which shows the statistical distribution of the developer-ChatGPT conversations we collected from GitHub based on the nine activities. Notably, Software Development is the most frequent activity where developers share ChatGPT conversations, accounting for a total of 1,059 (41.57%) interactions. Most of these shared developer-ChatGPT conversations are from Code (775, 30.43%), Commits (108, 4.2%), and Issues (107, 4.2%), underscoring the deep involvement and widespread use of ChatGPT in modifying and improving codebases. Software Maintenance and Evolution is the second largest development-related activity, with a total of 844 (33.14%) interactions. Note that, in this activity, there is a high volume of Commits (579, 22.73%) and a substantial number of Pull Requests (116, 4.55%), illustrating the ongoing need for developers to refine and optimize existing software leveraging ChatGPT. Compared to the above two dominant activities, Information Seeking is less frequent, yet still involving 172 (6.75%) developer-ChatGPT conversations. Moreover, Software Deployment mainly includes deployment problems in Code, followed by moderate developer-ChatGPT conversations in Commits and Issues. Software Design, Software Testing, and Data Handling show relatively balanced developers-ChatGPT conversations in Code, Commits, Issues, and Pull Requests. Software Management involves ChatGPT links mainly in Commits with Requirements Analysis having the least developer-ChatGPT conversations.
Figure 9 offers a different perspective on a visual distribution heat map, illustrating the development-related activities involving shared ChatGPT conversations across the five sources. Each data source is illustrated with its respective percentage allocation, offering insights into how ChatGPT is utilized across different sources on GitHub.
Code (1,105): The dominant activity where ChatGPT conversations are shared is Software Development, accounting for 70.14% of all activities, while the proportions of Software Maintenance and Evolution (6.06%) and Information Seeking (9.68%) are much lower. This implies that developers most frequently rely on ChatGPT to directly support core development tasks, such as generating and understanding code. Other activities, such as Software Deployment (4.62%), Software Design (3.35%), and Software Testing (3.44%), represent comparatively lower percentages, further reinforcing that most activities in Code on GitHub involve direct code generation rather than less prominent activities like design or testing.
Commits (823): This source reveals a strong focus on Software Maintenance and Evolution (70.35%), followed by Software Development (13.00%), emphasizing ChatGPT’s role in iterative improvement and code refinement through code modifications. This aligns with our intuition, given that GitHub Commits generally indicates that developers are actively working on improving and maintaining the software, committing updates regularly to ensure software stability and improvements. Other activities like Software Management (6.80%) and Software Deployment (4.25%) contribute less, indicating relatively limited engagement with ChatGPT for software management or deployment.
Issues (295): Software Development again ranks highest at 36.61%, followed by Software Maintenance and Evolution (25.08%), indicating ChatGPT’s utility in addressing core development needs like debugging and code refinement. Besides, Information Seeking (9.49%) is a notable activity, highlighting ChatGPT’s role as a tool for knowledge retrieval within collaborative problem-solving. Software Deployment (9.15%) and Software Design (7.80%) show lower percentages, suggesting that developers also turn to ChatGPT for deployment-related troubleshooting and refining design-level decisions during issue resolution.
Pull Requests (266): This source is primarily related to Software Maintenance and Evolution (43.61%), followed by Software Development (18.42%) and Data Handling (12.41%). This distribution indicates that developers often turn to ChatGPT for quality assurance, feature enhancements, and data processing. Other activities show less involvement in pull requests, further highlighting ChatGPT’s role as a supporting tool for ensuring software quality and consistency during code review.
Discussions (58): Software Development again emerges as the leading activity (34.48%), and followed by Information Seeking ranks second at 24.14%, illustrating ChatGPT’s value as an on-demand knowledge source. Furthermore, Software Deployment (15.52%) and Software Maintenance and Evolution (13.79%) follow, where ChatGPT helps to troubleshoot and guide deployment. Software Design (6.90%) also appears occasionally, showing that discussions often involve high-level architecture and planning with ChatGPT’s assistance.
4.4. RQ4: Tasks of ChatGPT Conversations Involved in Software Development
To answer RQ4, we investigated the specific tasks within developer-ChatGPT interactions during development. Figure 10 presents a Sankey diagram that visually breaks down nine development-related activities (see Section 4.3) into specific SE tasks with fine granularity. The Sankey diagram effectively illustrates the relationships between the five data sources, nine development-related activities, and 39 detailed SE tasks derived from ChatGPT interactions observed on our collected GitHub data, reflecting how developers utilize ChatGPT in their daily workflows.
According to Figure 10, the left five data sources not only show the diversity of available information but also underscore the multifaceted nature of collaborative software development, laying the foundation for subsequent activities. The middle nodes encompass a wide spectrum of development-related activities, from requirements analysis to software maintenance and evolution. The right part of Figure 10 breaks the nine development-related activities down into 39 specific SE tasks. In essence, the Sankey diagram visually traces the flow from each data source through the middle layer of development-related activities to specific SE tasks, revealing the deep integration of ChatGPT into the SDLC.
The most prominent flow originates from Code (1,105) and Software Development (1,059), channeling into Code Generation and Completion (888), which reflects the frequent practice of using ChatGPT’s capability to automate code generation during software development. Similarly, Commits (823) strongly correlate with Software Maintenance and Evolution (844), contributing to Code Modification and Optimization (686) and Code Review and Debugging (184), further reinforcing the focus on iterative code refinement. SE Tasks like Test Case Generation (59) derive from Software Testing (83), suggesting that ChatGPT aids in quality assurance and defect resolution.
Notably, tasks such as Deployment and Configuration (123) and Documentation Optimization (59) connect to Software Deployment (137) and Software Management (63), indicating auxiliary uses in deployment pipelines and project coordination. In contrast, underrepresented tasks like System Architecture Design (28) and Requirements Elicitation (5) reveal limited engagement of ChatGPT with specialized domains and early-phase development activities. This disparity also suggests developers primarily leverage ChatGPT for execution-phase tasks (e.g., code generation and completion) rather than tasks related to system design or requirements analysis (e.g., requirements elicitation).
Overall, beyond illustrating the quantitative distribution of interactions, Figure 10 offers valuable qualitative insights into the relational dynamics between data sources, development-related activities, and specific SE tasks. This visualization enables a nuanced understanding of how developers utilize ChatGPT across a broad spectrum of SE tasks, ranging from routine coding to complex problem-solving and architectural decisions. This integrated view can serve as a valuable framework for both academic research and practical applications in understanding the multifaceted nature of modern software engineering within collaborative environments.
5. Discussions and Implications
5.1. Interpretation of the Results
5.1.1. Characteristics of ChatGPT Usage on GitHub
According to the results of RQ1, the pronounced dominance of Code in shared ChatGPT interactions (43.4%) indicates ChatGPT’s prevalence in code-related tasks like generating code snippets and providing code explanations. Similarly, Commits (32.3%), holding the second largest proportion, reflects developers’ trust in ChatGPT to refine and validate code changes before integration. The sharp increase in ChatGPT’s usage following the release of its sharing feature (peaking around August 2023, just three months later after the release) illustrates the rapid adaptability of developers in embracing AI tools that can be smoothly integrated into their existing workflows. In contrast, the relatively limited engagement in Discussions suggests that developers tend to favor using ChatGPT for structured, task-specific interactions over open-ended conversations, aligning with the efficiency-focused demands of modern software development practices (Vaillant et al., 2024). Interestingly, the usage of ChatGPT’s sharing feature peaked in August 2023, after which the growth rate began to decline. This trend likely reflects an initial surge in user engagement driven by the novelty effect following the feature’s launch, which gradually tapered off as the novelty diminished.
Besides, the analysis of prompt turns distributions and descriptions regarding the shared ChatGPT links reveals a versatile integration of ChatGPT into the development process. The short and task-focused prompts (e.g., those prompts with only one or two turns) indicate that developers typically seek immediate and concise support for specific issues (e.g., concept interpretation, code modification), which is crucial in fast-paced coding environments. Moreover, the strong tendency to include contextual descriptions of shared ChatGPT links (especially in Commits, Pull Requests, and Discussions) reflects an underlying emphasis on transparency and clarity within collaborative development environments. This practice not only facilitates peer review and knowledge sharing, but also signifies a cultural shift toward embracing AI-assisted insights as an integral part of SDLC. As developers continue to employ GenAI tools like ChatGPT, these tools are well-positioned to enhance both individual efficiency and collective innovation in SE workflows (Vaillant et al., 2024).
5.1.2. Developers’ Purposes for Sharing ChatGPT Conversations
ChatGPT demonstrates notable proficiency in automating routine development tasks, as reflected in the dominance of Task Delegation within Commits (85.66%) and Pull Requests (88.35%). This trend points to a tangible reduction in developer workload for repetitive or procedural activities. However, ChatGPT’s relatively low engagement in Problem Resolution (11.42%) and Concept Interpretation (5.17%) highlights its current limitations in handling nuanced, context-dependent challenges, thereby reaffirming the indispensable role of human expertise in complex decision-making processes (e.g., trade-off analysis considering business context and quality concerns when making architectural decisions).
Conversely, ChatGPT exhibits promising potential as a brainstorming and ideation tool, particularly in Discussions and Issue, in which a higher engagement is observed for Knowledge Acquisition (20.69%) and Solution Recommendation (17.24%). ChatGPT’s markedly lower usage in code-centric tasks, such as Commits (0.61%), underscores its limited reliability in scenarios demanding exactness and accountability.
The observed divergence of developers’ purposes for using ChatGPT across different data sources highlights the need for future GenAI tools to be context-aware and provide tailored support aligned with specific stages of SDLC. For instance, integrating ChatGPT into issue-tracking systems could augment early-stage ideation and collaborative problem-solving, while ChatGPT’s automation capabilities could be beneficial for improving pull request workflows.
5.1.3. Development-Related Activities Involving Sharing ChatGPT Conversations
The results of RQ3 reveal ChatGPT’s dual role in supporting both SDLC activities (e.g., Software Design) and emerging AI-augmented activities like Information Seeking and Data Handling. The prominence of Software Development (in Code) and Software Maintenance and Evolution (in Commits) highlights ChatGPT’s value in accelerating the coding process and refining existing functionality. However, ChatGPT’s presence across diverse data sources on GitHub, for example, aiding problem-solving in Issues, quality assurance in Pull Requests, and knowledge retrieval in Discussions, demonstrates ChatGPT’s adaptability to context-specific activities. This suggests that ChatGPT is not merely a code generator but a versatile assistant integrated throughout the SDLC phases to streamline SE workflows.
Compared to the predominantly code-centric activities mentioned above, the relatively lower involvement of ChatGPT in Requirements Analysis and Software Management likely reflects ChatGPT’s intrinsic limitations in these two activities because domain knowledge and project-specific information are still needed. As of yet, there is no doubt that such areas still rely fundamentally on human expertise and cannot be fully replaced by GenAI. Additionally, the emergence of Information Seeking as a distinct activity indicates a paradigm shift: developers increasingly treat GenAI tools as a knowledge partner that integrates problem-solving and pair programming. This challenges traditional SDLC taxonomies, promoting a redefinition of development-related activities to account for hybrid human-AI collaboration.
5.1.4. SE Tasks Involving ChatGPT’s Engagement in Software Development
The findings for RQ4 indicate that ChatGPT is primarily employed for the automation of coding-related tasks, specifically Code Generation & Completion (888 instances) and Code Modification & Optimization (686 instances), underscoring its value in supporting coding and maintenance tasks. In contrast, early-phase tasks like Requirements Elicitation (5 instances) and System Architecture Design (28 instances) exhibit minimal usage of ChatGPT. One possible reason is that developers found ChatGPT unable to fulfill certain tasks owing to its limitations in understanding complex requirements and contributing to system design (Liang et al., 2024; Vaillant et al., 2024; Fan et al., 2023; Hou et al., 2023). Moreover, the distribution regarding the SE tasks in Figure 10 could also confirm such a tendency: developers are predominantly using ChatGPT for on-demand assistance (e.g., coding, maintenance, troubleshooting tasks), rather than as a collaborator for architecture-level planning (e.g., architectural design and optimization). This observation raises a question about whether ChatGPT’s limitations in contextual reasoning (Bang et al., 2023) and domain-specific knowledge constrain its applications in early development stages (Cong-Lem et al., 2025). Consequently, tasks that require intensive human judgment and holistic system understanding (e.g., Requirements Elicitation), remain underexplored in LLM-assisted SE workflows.
In addition, ChatGPT’s relatively low engagement with Software Testing (e.g., Test Case Generation) points to a potential gap in the adoption of LLM-powered testing automation (Li et al., 2023). Although LLM-generated tests could reduce manual effort, developers may distrust automatically-generated tests to accurately model complex scenarios, thereby preferring human oversight (Ouédraogo et al., 2024).
5.2. Implications
Establish New Paradigms for Software Development.
Our study results reveal how ChatGPT is being applied on GitHub, yielding usage characteristics and patterns that can help tool vendors and platform designers tailor LLM-powered features that better support developers in their daily work. Therefore, the integration of LLM-driven assistance should not merely be seen as an incremental improvement but as a fundamental shift in the SE paradigms, shaping the future of SE research, such as both development processes and decision-making in the field (Gao et al., 2025).
For practitioners and tool vendors, the usage of ChatGPT in SE tasks shows a strong focus on immediate assistance provided by ChatGPT for coding and maintenance tasks. There is significant potential for deep integration of GenAI tools like ChatGPT into developer ecosystems (e.g., IDEs, version control) to enhance context-aware support, which dynamically adapts to the developer’s current code (e.g., automatically searching relevant API examples based on code context during pull-request reviews). Meanwhile, proposing guidelines and processes for reviewing and validating code or content generated by ChatGPT can further safeguard code quality and reliability.
For researchers, the findings point to promising opportunities to extend ChatGPT’s utility in certain areas, such as requirements engineering, by tailoring LLMs to domain-specific SE workflows. For example, the concentration of SE tasks in Software Development and Software Maintenance & Evolution implies that ChatGPT is primarily perceived as a coding assistant. Moreover, the findings highlight research opportunities to integrate ChatGPT in certain tasks (e.g., System Architecture Design), thereby improving its capability to boost productivity in underexploited areas. Besides, the study underscores ChatGPT’s transformative potential - not only in reshaping how developers engage in SE tasks, but also in redefining what constitutes a development-related activity in the AI era. This evolution calls for foundational knowledge in software engineering, such as SWEBOK (Washizaki et al., 2024), to adapt alongside emerging GenAI-driven software development practices.
Nevertheless, facing the rapid advances in the field of intelligent software engineering (a.k.a AI4SE) (Hou et al., 2023; Gao et al., 2025), our SE research community should develop new methods, standards, and ethical guidelines that integrate intelligent AI tools or components into the SDLC, to ensure that human-AI collaboration improves code quality, developer productivity, and system maintainability (Gao et al., 2025; Terragni et al., 2024).
Assess the Impact of ChatGPT on Developer Productivity and Code Quality.
While this study provides valuable insights regarding the daily use of ChatGPT during software development, future research could delve deeper into the impact of ChatGPT on developer productivity. This could involve conducting controlled experiments to compare outcomes with and without ChatGPT’s assistance.
For practitioners, understanding the impact of ChatGPT is crucial for strategic technology adoption during development. Although GenAI tools like ChatGPT exhibit great potential in enhancing developer efficiency (Vaillant et al., 2024), their limitations, such as the propensity to generate inaccurate or misleading information (often referred to as hallucination (Achiam et al., 2023)) or even unsafe code content (Chen et al., 2025), cannot be overlooked. Therefore, for researchers, this necessitates the collection of robust empirical evidence to assess the reliability and practical utility of ChatGPT in SE tasks. Future investigations should focus on determining the contexts in which these tools deliver substantial benefits and the contexts where they may introduce risks or inefficiencies (Yu et al., 2024). Addressing these concerns will be critical to ensuring the responsible and effective integration of LLMs into modern software development workflows.
Stay Alert on the Risks of LLM-Generated Content.
Our results reveal that developers utilize ChatGPT in various development-related activities (particularly Software Development and Software Maintenance & Evolution). However, errors in ChatGPT’s outputs could potentially introduce inefficiencies or vulnerabilities in software systems (Liu et al., 2024b, a). Moreover, excessive cognitive reliance on GenAI tools might hinder the development of essential manual problem-solving skills, especially for junior developers. Given the potential of errors and biases in LLM-generated content (Chen et al., 2025; Li et al., 2024), it is imperative to establish robust safeguards, such as human oversight, rigorous validation processes, and strict adherence to ethical and legal standards. These measures are critical to maintaining the quality, security, and trustworthiness of AI-assisted development.
Specifically, it is essential to investigate the potential risks and limitations of using ChatGPT during software development. This includes assessing concerns related to intellectual property, fairness, and the introduction of biases into LLM-generated content. Practitioners must adopt strategies to mitigate these risks, such as incorporating human inspection for critical tasks, ensuring transparency in the LLM generation process since humans, not AI, must bear responsibility for the consequences of incorrectly generated content, and rigorously following ethical guidelines for using GenAI in software development (Akbar et al., 2023). Continuous evaluation of GenAI tools and their outputs is crucial to ensure the trustworthiness and reliability of their generated results.
6. Threats to Validity
In this section, we discuss the potential threats to the validity related to our study and the adopted measures mitigating these threats according to the guidelines of Wohlin et al. (Wohlin et al., 2012). Internal validity is not considered in this work, as no causal relationships between variables are addressed.
Construct validity: Potential threats of this study lie in the mined ChatGPT links and their content that developers shared on GitHub. Not all shared ChatGPT links reflect meaningful interactions, and some may be noise, off-topic, or duplicates. For example, many shared ChatGPT links are not relevant to software development, or certain development-related ChatGPT links are duplicated from different sources (e.g., one link occurs in both Code and Issues). To mitigate these threats, we employed rigorous data cleaning techniques to eliminate noise data (e.g., irrelevant and duplicates) and ensure the uniqueness of mined data through a specific ID filtering process.
External validity: A relevant threat to the generalizability of our findings concerns the temporal stability of the observed findings. Our analysis focuses on developer-ChatGPT interactions on GitHub from May 2023 to June 17, 2024, a period that coincides with the release of ChatGPT’s sharing feature. As GenAI models and their applications continue to evolve, the patterns of usage and perceived effectiveness may also shift over time. Additionally, the findings may not fully generalize to other open-source platforms or proprietary development environments, which may involve different user demographics or AI tools. While our study captures a significant portion of ChatGPT usage among developers, it may not fully represent the broader spectrum of usage across all development contexts. Nevertheless, given GitHub’s prevalence as a leading platform for collaborative software development, we believe that our findings reasonably reflect broader trends within the developer community.
Reliability: The threats to the replicability of our study primarily stem from the process of data collection, cleaning, and labeling. To alleviate these threats, we have provided a detailed description of our study design in Section 3, outlining specific steps involved in data curation. Specifically, we conducted a pilot data labeling exercise to ensure a consistent understanding of labeling criteria among authors. Any discrepancies and disagreements were resolved through discussions. Furthermore, to promote transparency and facilitate reproducibility, we have made the datasets publicly available online (Li et al., 2025). This enables researchers to independently replicate our findings and conduct further research based on our shared data.
7. Conclusions and Future Work
The integration of LLMs, particularly ChatGPT, into software development marks a significant milestone in the evolution of SE practices. We are witnessing a paradigm shift from traditional development to AI-assisted development, streamlining SE workflows and reducing development time and effort. This study provides a comprehensive analysis of ChatGPT’s practical usability and its impact on AI-assisted software development through the curated DevChat dataset, comprising 2,547 unique shared ChatGPT links from GitHub.
Our findings reveal the characteristics of ChatGPT’s usage on GitHub, such as ChatGPT’s prompt turns distribution, and accompanying description. Moreover, the results uncover five primary purposes for which developers share developer-ChatGPT conversations, with Task Delegation being the dominant purpose, primarily to automate work. Furthermore, our results emphasize the widespread adoption of ChatGPT across nine development-related activities, with Code and Commits serving as the main data sources, and Software Development, Software Maintenance and Evolution as the dominant activities in which developers frequently utilize ChatGPT. In addition, we established a comprehensive mapping framework to analyze the specific SE tasks for which developers employ ChatGPT during their daily routines. Overall, our study demonstrates the pivotal role of GenAI tools represented by ChatGPT in supporting a wide range of SE activities across the SDLC.
The results contribute to a comprehensive understanding of the growing impact of ChatGPT on SE practices. The results also provide actionable insights for tool builders aiming to enhance the usability of AI-assisted development tools and for researchers studying the implications of AI adoption in SE workflows. Future work could consider extending the analysis beyond the temporal scope of the DevChat dataset to examine how ChatGPT usage evolves alongside ongoing advancements in LLMs. Furthermore, researchers could study the impacts of developers’ prompts on ChatGPT’s outputs, and conduct comparative analyses between ChatGPT and other emerging LLMs (e.g., DeepSeek R1, Llama 3.3) to better understand their respective strengths and weaknesses. Ultimately, continued exploration of ChatGPT’s role in SE will be instrumental in shaping the future of AI-driven software development practices.
Data Availability
We provide our collected, extracted, and processed datasets online at (Li et al., 2025).
Acknowledgements.
This work has been partially supported by the National Natural Science Foundation of China (NSFC) with Grant Nos. 62402348 and 62172311, and the Major Science and Technology Project of Hubei Province under Grant No. 2024BAA008.References
- (1)
- Achiam et al. (2023) Josh Achiam, Steven Adler, Sandhini Agarwal, Lama Ahmad, Ilge Akkaya, Florencia Leoni Aleman, Diogo Almeida, Janko Altenschmidt, Sam Altman, and Shyamal Anadkat. 2023. Gpt-4 technical report. Technical Report.
- Aguiar et al. (2024) Lucas Aguiar, Matheus Paixao, Rafael Carmo, Edson Soares, Antonio Leal, Matheus Freitas, and Eliakim Gama. 2024. Multi-language Software Development in the LLM Era: Insights from Practitioners’ Conversations with ChatGPT. In Proceedings of the 18th ACM/IEEE International Symposium on Empirical Software Engineering and Measurement (ESEM). ACM, 489–495.
- Akbar et al. (2023) Muhammad Azeem Akbar, Arif Ali Khan, and Peng Liang. 2023. Ethical aspects of ChatGPT in software engineering research. IEEE Transactions on Artificial Intelligence 6, 2 (2023), 254–267.
- Anthropic, Claude (2025) Anthropic, Claude. 2025. Retrieved March 20, 2025 from https://claude.ai/chats.
- Bang et al. (2023) Yejin Bang, Samuel Cahyawijaya, Nayeon Lee, Wenliang Dai, Dan Su, Bryan Wilie, Holy Lovenia, Ziwei Ji, Tiezheng Yu, Willy Chung, et al. 2023. A multitask, multilingual, multimodal evaluation of ChatGPT on reasoning, hallucination, and interactivity. In Proceedings of the 13th International Joint Conference on Natural Language Processing and the 3rd Conference of the Asia-Pacific Chapter of the Association for Computational Linguistics (IJCNLP). ACL, 675–718.
- Charmaz (2014) Kathy Charmaz. 2014. Constructing Grounded Theory. Sage.
- ChatGPT Shared Links FAQ (2025) ChatGPT Shared Links FAQ. 2025. Retrieved March 20, 2025 from https://help.openai.com/en/articles/7925741-chatgpt-shared-links-faq.
- Chen et al. (2025) Yuchen Chen, Weisong Sun, Chunrong Fang, Zhenpeng Chen, Yifei Ge, Tingxu Han, Quanjun Zhang, Yang Liu, Zhenyu Chen, and Baowen Xu. 2025. Security of Language Models for Code: A Systematic Literature Review. ACM Transactions on Software Engineering and Methodology (2025), 1–66.
- Cohen (1960) Jacob Cohen. 1960. A coefficient of agreement for nominal scales. Educational and Psychological Measurement 20, 1 (1960), 37–46.
- Cong-Lem et al. (2025) Ngo Cong-Lem, Ali Soyoof, and Diki Tsering. 2025. A systematic review of the limitations and associated opportunities of ChatGPT. International Journal of Human–Computer Interaction 41, 7 (2025), 3851–3866.
- Fan et al. (2023) Angela Fan, Beliz Gokkaya, Mark Harman, Mitya Lyubarskiy, Shubho Sengupta, Shin Yoo, and Jie M Zhang. 2023. Large language models for software engineering: Survey and open problems. In Proceedings of the IEEE/ACM International Conference on Software Engineering: Future of Software Engineering (ICSE-FoSE). IEEE, 31–53.
- Gao et al. (2025) Cuiyun Gao, Xing Hu, Shan Gao, Xin Xia, and Zhi Jin. 2025. The Current Challenges of Software Engineering in the Era of Large Language Models. ACM Transactions on Software Engineering and Methodology (2025).
- GitHub, Copilot (2025) GitHub, Copilot. 2025. Retrieved March 20, 2025 from https://github.com/features/copilot.
- Github Language Stats (2025) Github Language Stats. 2025. Retrieved March 20, 2025 from https://madnight.github.io/githut/#/pull_requests/2024/1.
- Github REST API documentation (2022) Github REST API documentation. 2022. Retrieved March 20, 2025 from https://docs.github.com/en/rest?apiVersion=2022-11-28.
- Google, Gemini (2025) Google, Gemini. 2025. Retrieved March 20, 2025 from https://gemini.google.com/.
- Guo et al. (2024) Qi Guo, Junming Cao, Xiaofei Xie, Shangqing Liu, Xiaohong Li, Bihuan Chen, and Xin Peng. 2024. Exploring the potential of ChatGPT in automated code refinement: An empirical study. In Proceedings of the 46th IEEE/ACM International Conference on Software Engineering (ICSE). ACM, 1–13.
- Hao et al. (2024) Huizi Hao, Kazi Amit Hasan, Hong Qin, Marcos Macedo, Yuan Tian, Steven HH Ding, and Ahmed E Hassan. 2024. An Empirical Study on Developers Shared Conversations with ChatGPT in GitHub Pull Requests and Issues. Empirical Software Engineering 29, 6 (2024), 150.
- Hou et al. (2023) Xinyi Hou, Yanjie Zhao, Yue Liu, Zhou Yang, Kailong Wang, Li Li, Xiapu Luo, David Lo, John Grundy, and Haoyu Wang. 2023. Large language models for software engineering: A systematic literature review. ACM Transactions on Software Engineering and Methodology 33, 8 (2023), 1–79.
- Jin et al. (2023) Matthew Jin, Syed Shahriar, Michele Tufano, Xin Shi, Shuai Lu, Neel Sundaresan, and Alexey Svyatkovskiy. 2023. Inferfix: End-to-end program repair with llms. In Proceedings of the 31st ACM Joint European Software Engineering Conference and Symposium on the Foundations of Software Engineering (ESEC/FSE). ACM, 1646–1656.
- Kabir et al. (2024) Samia Kabir, David N Udo-Imeh, Bonan Kou, and Tianyi Zhang. 2024. Is Stack Overflow obsolete? An empirical study of the characteristics of ChatGPT answers to Stack Overflow questions. In Proceedings of the 44th CHI Conference on Human Factors in Computing Systems (CHI). ACM, 1–17.
- Landis and Koch (1977) J Richard Landis and Gary G Koch. 1977. The measurement of observer agreement for categorical data. Biometrics (1977), 159–174.
- Li et al. (2025) Ruiyin Li, Peng Liang, Yifei Wang, Yangxiao Cai, Weisong Sun, and Zengyang Li. 2025. Dataset for the Paper: Unveiling the Role of ChatGPT in Software Development: Insights from Developer-ChatGPT Interactions on GitHub.
- Li et al. (2023) Tsz-On Li, Wenxi Zong, Yibo Wang, Haoye Tian, Ying Wang, Shing-Chi Cheung, and Jeff Kramer. 2023. Nuances are the Key: Unlocking ChatGPT to Find Failure-Inducing Tests with Differential Prompting. In Proceedings of the 38th IEEE/ACM International Conference on Automated Software Engineering (ASE). IEEE, 14–26.
- Li et al. (2024) Xinyue Li, Zhenpeng Chen, Jie M. Zhang, Yiling Lou, Tianlin Li, Weisong Sun, Yang Liu, and Xuanzhe Liu. 2024. Benchmarking Bias in Large Language Models during Role-Playing. arXiv preprint arXiv:2411.00585 (2024).
- Liang et al. (2024) Jenny T Liang, Chenyang Yang, and Brad A Myers. 2024. A large-scale survey on the usability of AI programming assistants: Successes and challenges. In Proceedings of the 46th IEEE/ACM International Conference on Software Engineering (ICSE). ACM, 1–13.
- Liu et al. (2024b) Jiawei Liu, Chunqiu Steven Xia, Yuyao Wang, and Lingming Zhang. 2024b. Is your code generated by ChatGPT really correct? Rigorous evaluation of large language models for code generation. Advances in Neural Information Processing Systems 36 (2024), 21558–21572.
- Liu et al. (2024a) Zhijie Liu, Yutian Tang, Xiapu Luo, Yuming Zhou, and Liang Feng Zhang. 2024a. No need to lift a finger anymore? Assessing the quality of code generation by ChatGPT. IEEE Transactions on Software Engineering 50, 6 (2024), 1548–1584.
- Lutellier et al. (2020) Thibaud Lutellier, Hung Viet Pham, Lawrence Pang, Yitong Li, Moshi Wei, and Lin Tan. 2020. Coconut: combining context-aware neural translation models using ensemble for program repair. In Proceedings of the 29th ACM SIGSOFT international symposium on software testing and analysis (ISSTA). ACM, 101–114.
- Meta, Llama (2025) Meta, Llama. 2025. Retrieved March 20, 2025 from https://www.llama.com/.
- OpenAI, ChatGPT (2025) OpenAI, ChatGPT. 2025. Retrieved March 20, 2025 from https://www.openai.com/.
- Ouédraogo et al. (2024) Wendkûuni C Ouédraogo, Yinghua Li, Kader Kaboré, Xunzhu Tang, Anil Koyuncu, Jacques Klein, David Lo, and Tegawendé F Bissyandé. 2024. Test smells in LLM-Generated Unit Tests. arXiv preprint arXiv:2410.10628 (2024).
- Poldrack et al. (2023) Russell A Poldrack, Thomas Lu, and Gašper Beguš. 2023. AI-assisted coding: Experiments with GPT-4. arXiv preprint arXiv:2304.13187 (2023).
- Prenner et al. (2022) Julian Aron Prenner, Hlib Babii, and Romain Robbes. 2022. Can OpenAI’s codex fix bugs? An evaluation on QuixBugs. In Proceedings of the 3rd International Workshop on Automated Program Repair (APR). IEEE, 69–75.
- REST API endpoints for search (2025) REST API endpoints for search. 2025. Retrieved March 20, 2025 from https://docs.github.com/en/rest/search/search?apiVersion=2022-11-28#about-search.
- Sobania et al. (2023) Dominik Sobania, Martin Briesch, Carol Hanna, and Justyna Petke. 2023. An analysis of the automatic bug fixing performance of ChatGPT. In The 4th International Workshop on Automated Program Repair (APR@ICSE). IEEE/ACM, 23–30.
- Stol et al. (2016) Klaas-Jan Stol, Paul Ralph, and Brian Fitzgerald. 2016. Grounded theory in software engineering research: A critical review and guidelines. In Proceedings of the 38th International Conference on Software Engineering (ICSE). ACM, 120–131.
- Su and McMillan (2024) Chia-Yi Su and Collin McMillan. 2024. Distilled GPT for source code summarization. Automated Software Engineering 31, 1 (2024), 22.
- Sun et al. (2025) Weisong Sun, Yun Miao, Yuekang Li, Hongyu Zhang, Chunrong Fang, Yi Liu, Gelei Deng, Yang Liu, and Zhenyu Chen. 2025. Source Code Summarization in the Era of Large Language Models. In Proceedings of the 47th International Conference on Software Engineering (ICSE). IEEE, 419–431.
- Terragni et al. (2024) Valerio Terragni, Partha Roop, and Kelly Blincoe. 2024. The Future of Software Engineering in an AI-Driven World. In Proceedings of the Workshop on Software Engineering in 2030 (SE 2030). ACM, 1–6.
- Vaillant et al. (2024) Thiago S Vaillant, Felipe Deveza de Almeida, Paulo Anselmo Neto, Cuiyun Gao, Jan Bosch, and Eduardo Santana de Almeida. 2024. Developers’ Perceptions on the Impact of ChatGPT in Software Development: A Survey. arXiv preprint arXiv:2405.12195 (2024).
- Vaithilingam et al. (2022) Priyan Vaithilingam, Tianyi Zhang, and Elena L Glassman. 2022. Expectation vs. experience: Evaluating the usability of code generation tools powered by large language models. In Proceedings of the 42nd CHI Conference on Human Factors in Computing Systems (CHI): Extended Abstracts. ACM, 1–7.
- Vaswani et al. (2017) Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones, Aidan N Gomez, Lukasz Kaiser, and Illia Polosukhin. 2017. Attention is all you need. In Proceedings of the 31st Annual Conference on Neural Information Processing Systems (NIPS). NeurIPS Proceedings, 5998–6008.
- Washizaki et al. (2024) Hironori Washizaki et al. 2024. Guide to the Software Engineering Body of Knowledge (SWEBOK (R)): Version 4.0. IEEE.
- Wessel et al. (2023) Mairieli Wessel, Joseph Vargovich, Marco A Gerosa, and Christoph Treude. 2023. Github Actions: The Impact on the Pull Request Process. Empirical Software Engineering 28, 6 (2023), 131.
- Wohlin et al. (2012) Claes Wohlin, Per Runeson, Martin Höst, Magnus C Ohlsson, Björn Regnell, and Anders Wesslén. 2012. Experimentation in Software Engineering. Springer Science & Business Media.
- Xiao et al. (2024) Tao Xiao, Christoph Treude, Hideaki Hata, and Kenichi Matsumoto. 2024. DevGPT: Studying Developer-ChatGPT Conversations. In Proceedings of the 21st International Conference on Mining Software Repositories (MSR). ACM, 227–230.
- Yu et al. (2024) Xiao Yu, Lei Liu, Xing Hu, Jacky Wai Keung, Jin Liu, and Xin Xia. 2024. Fight Fire with Fire: How Much Can We Trust ChatGPT on Source Code-Related Tasks? IEEE Transactions on Software Engineering 50, 12 (2024), 3435–3453.
- Zhang et al. (2023) Quanjun Zhang, Tongke Zhang, Juan Zhai, Chunrong Fang, Bowen Yu, Weisong Sun, and Zhenyu Chen. 2023. A Critical Review of Large Language Model on Software Engineering: An Example from ChatGPT and Automated Program Repair. arXiv preprint arXiv:2310.08879 (2023).